### v1

# Data Science Education in Undergraduate Physics: Lessons Learned from a Community of Practice

Abstract
With the increasing availability of diverse datasets, ranging from small-scale experimental data points to large and complex data repositories and powerful data analysis tools, it is increasingly important that physics educators equip their students with the skills to work with data effectively. However, many educators may lack the necessary training and expertise in data science to teach these skills. To address this gap, we created the Data Science Education Community of Practice (DSECOP), bringing together graduate students and physics educators from different institutions and backgrounds to share best practices and lessons learned in integrating data science into undergraduate physics education. In this article, we present insights and experiences from this community of practice, highlighting key strategies and challenges in incorporating data science into the introductory physics curriculum. Our goal is to provide guidance and inspiration to educators who seek to integrate data science into their teaching, helping to prepare the next generation of physicists for a data-driven world.
I Introduction
Data has always been an essential component of the scientific method. From Kepler’s deriving laws of planetary motion by analyzing Brahe’s astronomical observations to the detection of the Higgs boson, data has played a crucial role in advancing our understanding of the natural world. Traditionally, data analysis in physics primarily revolved around structured datasets, focusing on applying theoretical concepts to data derived from experiments or simulations. In the current age of big data, physicists have access to an unprecedented volume and variety of data, enabling new discoveries that would have been impossible without advanced data analysis techniques. This involves not just analytical skills but also proficiency in advanced data management, statistical modeling, and computational techniques. In light of these developments, it is essential that physics students learn the skills to work with data effectively.
However, data science is not traditionally taught in undergraduate physics classes. The historical focus on theoretical and experimental physics leaves little room for data science education with physics applications. Additionally, many physics educators lack the training and expertise to teach data science effectively, in particular, because of the fast-paced evolution of tools and techniques. As illustrated in figure 1, the importance of data science in physics research and industry has been growing rapidly. There is a need to integrate data science into undergraduate physics curricula to prepare students for the changing demands of the future.
Figure 1: Percentage of academic papers per year uploaded to arXiv’s physics section that mention “machine learning” anywhere in the article. The red dashed line represents the inception of the American Physical Society’s Topical Group on Data Science (GDS).
It is difficult to define data science comprehensively, partly because many definitions tend to emphasize the specific activities of practitioners within their respective fields provost2013data . However, we will provide an operational definition. The term was first used in the 1960s in military and corporate contexts alvarado_data_2023 . While data science is an umbrella term Irizarry2020Role , the NIST Big Data Interoperability Framework nist_big_data_public_working_group_definitions_and_taxonomies_subgroup_nist_2015 provides a succinct definition covering various aspects of data science:
“Data science is the extraction of actionable knowledge directly from data through a process of discovery, or hypothesis formulation and hypothesis testing.”
Data science uses statistical and computational methods to extract insights from data. It includes data collection, cleaning, preparation, exploratory data analysis, statistical modeling and inference, machine learning (ML), and data visualization. It has been used to address questions in various domains, such as biology, education, physics, business, linguistics, or medicine oliver2021undergraduate .
Data science focuses on analyzing large and complex datasets to identify patterns and trends, gain insights, and make probabilistic predictions. It deals directly with real-world data. There are, however, implicit models in data science as well (with many more parameters), and we see that as ML models get closer to real-world relationships, their effectiveness improves. ML models already serve as faster and more generalizable surrogate models, replacing theory in a variety of domains carleoMachineLearningPhysical2019a ; karniadakisPhysicsinformedMachineLearning2021b . However, a key limitation of many current ML models is their lack of interpretability, often functioning as ‘black boxes’ that are not readily explainable. In the future, ML models may represent reality with interpretable models as theoretical physics does. While we are far from that, it is an exciting research direction for physicists.
Why is data science important, particularly for physics undergraduates?
  1. 1.
    Physics research is increasingly data-intensive (Data science for Physics): Many fields of physics, such as astronomy, particle physics, condensed matter physics, and biophysics, produce vast amounts of data that require sophisticated data analysis. In addition to theory, experiment, and simulations, data science/machine learning is the “fourth pillar” of science and has a place in the scientific method hey2009the . With the rise of massive data-intensive experiments and simulations, it has become essential that physicists employ statistical and machine learning techniques in their workflow. Proficient data science skills are needed to write programs for sifting through and generating insights from experiments and large-scale simulations that generate terabytes of data. Additionally, these skills can enhance the efficiency of experiments and simulations, allowing for fewer measurements while still answering physically meaningful questions. Data science is further utilized in collider physics hajer2020novelty , quantum physics das2019machine , in the search for new materials, as well as controlling dynamic systems degraveMagneticControlTokamak2022 .
  2. 2.
    The increasing importance of data science in the workforce (Data science for Future Careers): In the era of Industry 4.0schwabFourthIndustrialRevolution2017 , where automation, data exchange, and interconnectedness shape our global industries, data science skills have become even more paramount. Physics undergraduates equipped with these skills are well-positioned to navigate the challenges and opportunities of this new industrial landscape. Acknowledging the already packed physics curriculum, it’s essential to note that the incorporation of data science doesn’t detract from core physics topics. Instead, it provides an enriched framework to understand and apply these principles using contemporary data science techniques. As industries pivot towards data-driven decision-making, grounding physics students in data science amplifies their analytical prowess and broadens their career horizons.
  3. 3.
    Data science can enhance physics education (Data science for Education): Data science offers new and innovative ways to teach physics concepts and engage students in active learning by providing students with hands-on experience with real-world data sets. Due to curriculum pressure, it is difficult to target parts of physics courses to be replaced by data science education. Introducing data science should not take away from the fundamental physics curriculum but, on the contrary, should make it easier to teach and demonstrate the fundamental ideas using modern data science tools. More and more resources and tools are becoming available for teaching data science. Free and open-source software, online tutorials, and educational materials make it easier than ever to introduce data science in undergraduate physics education.
    The importance of data science and machine learning in education, particularly in the field of high-energy physics (HEP), has been clearly spelled out by Benelli et al. benelli2022data .
  4. 4.
    Growing overlap between physical concepts and machine learning techniques (Physics for data science): Recent research has demonstrated that physical concepts, such as diffusion, symmetry, and hyperbolic (relativistic) geometry, can be used to develop more powerful and efficient machine learning algorithms and data analysis techniques (examples, references). By exploring the connections between physics and data science, students can gain insights into the fundamental principles that underlie both fields and develop a more holistic view of the role of data in scientific discovery and innovation. The relationship between data science and physics is bidirectional. While data science provides powerful tools for analyzing complex physical phenomena, physics, in turn, provides a rich context for developing and refining data science methodologies carleoMachineLearningPhysical2019a . Physics-informed machine learning is a rapidly developing field combining black-box ML models with physics constraints karniadakisPhysicsinformedMachineLearning2021b .
Why is it important for physics professors to teach it (as opposed to computer scientists)? The report of the National Academies of Sciences, Engineering & Medicine (NASEM) on data science for undergraduates emphasizes developing data expertise through a curriculum that includes important data science concepts, applications to real-world problems with an understanding of limitations, and ethical concerns NAP25104 .
Existing undergraduate data science programs focus primarily on theoretical foundations and quantitative skills with very little domain knowledge outside of computer science oliver2021undergraduate . Graduates of these programs may lack the appropriate context for designing and evaluating domain-specific data science applications. The NASEM framework emphasizes the importance of domain knowledge for effectively applying data science. Substantial domain knowledge in data science training is needed Berthold2019What ; Irizarry2020Role .
Survey Results Suggest Data Science is largely not part of the undergraduate physics curriculum
In March 2022, we ran two surveys to better understand the data science and undergraduate physics education landscape. The first survey was for faculty to learn whether they were teaching data science in their undergraduate physics courses and, if so, in which courses and what data science skills. The second survey was for data scientists in the industry who are involved with hiring undergraduate physicists, with the goal of understanding what data science skills and knowledge they would expect a recent bachelor’s in physics to have for data science jobs. The surveys were based on the DSECOP leadership’s experience in data science and physics, as well as a previous local survey losert2020 . Each survey was pilot-tested to ensure coherence and inclusion of important questions.
We do not have a response rate as the surveys were sent through American Physical Society (APS) listservs (e.g., Forum on Education [FEd], GDS), message boards, and our various professional networks. We assume the response rate was not high, given that the APS GDS list-serv alone has over 2000 members. This is potentially due to the limited bandwidth many experienced during the Covid-19 pandemic. There is still valuable information to be shared as the responses have some diversity around institutions for faculty and companies for industry professionals.
Faculty Survey
We had 100 physics faculty responses from 78 US institutions (5 US institutions had two respondents), 7 international institutions, and 10 unknown. 30 respondents said they had taught data science in undergraduate physics courses in the past five years; 25 are from US institutions, and 5 are from international institutions.
8% indicated they teach data science in their introductory and intermediate/advanced physics courses. Throughout this manuscript, both intermediate and advanced physics courses will be referred to as “advanced” for brevity. 2% indicated they teach data science only in introductory physics. 5% indicated that their introductory physics course that teaches data science is focused on data science (e.g., Data Science for Physicists). 19% indicated they teach data science in only intermediate or advanced physics. 12% indicated that their advanced physics course is focused on data science.
We do not know if the respondents teach both introductory and advanced courses, thus the opportunity may not arise to teach data science in those courses. However, this does suggest that data science is more commonly taught in advanced undergraduate physics courses. Thus, we focus on advanced undergraduate physics courses in this section of the manuscript.
For advanced courses, data science is taught via in-class activities (N = 8), projects (N = 25), and homework exercises (N = 19). Some respondents include data science questions on quizzes or exams (N = 8).
Figure 2: Results of data skills survey for faculty. Relevant skills are listed on the x axis, with the different textures showing the proportion of responses received.
Figure 2 is a plot of the data skills and the statement that the respondents selected (N = 26). The majority of respondents (over 50%) selected “I teach this in my intermediate/advanced physics courses that use data science and I do assess students’ knowledge skills, or abilities in this area” for Experimental design and scientific method (e.g., designing experiments, data collection), Data Manipulation (e.g., transforming data, gaining insights from data), Statistics (e.g., model selection and uncertainty, A/B testing), and Data Visualization and Communication (e.g., plotting data, technical writing). Most selected “I do not teach this in my intermediate/advanced physics courses that use data science and have no plans to do so” for Data Management /Storage (e.g., data typologies and storage/cloud architectures).
The majority of respondents (N = 70) have not taught data science in their physics courses in the past 5 years. Out of the 70 respondents, 69 answered the question regarding their interest in teaching data science in their physics courses. 40.6% selected they would be interested, 40.6% were maybe, and 18.8% were not interested. 61 of the respondents answered the free-response question regarding why they do not teach data science in their physics courses. These responses were thematically coded. The respondents often gave multiple reasons. The top themes that emerged are in the table 1 below.
Theme Count Example quotes
Not relevant or necessary courses taught
20
“… I have not judged it to be a priority (or even relevant) in courses I am teaching on topics such as quantum mechanics, special relativity, and ‘honors’ level mechanics.”
“Students need to learn the conceptual framework of physics first. There is little point in teaching them to compute if they have nothing to compute.”
Challenges in changing or adding to an already full curriculum
18
“Courses are already full. Hard to decide what to drop to make room”
“No space in the curriculum. Need to focus on teaching students physics concepts and methods, as well as some basic computational skills.”
“…not sure how to incorporate since we don’t have room in the program for a separate course ”
Faculty do not know how to use data science
13
“I don’t understand machine learning well enough to teach it. Although, it seems no one does.”
“No background in big data, ML, or AI. ”
Faculty does not know how to include data in their physics courses
13
“I don’t know how I would integrate it into courses I have taught…”
“I have no direct experience, wouldn’t know where to start or what courses to add it to.”
Table 1: Thematic coding of faculty response regarding why they do not include data science in their undergraduate physics courses
Most respondents did not see data science as relevant or necessary to the courses they teach. These respondents were not opposed to data science; some used data science and AI in their research, but they felt that teaching such skills was part of the core curriculum or could be learned elsewhere (e.g., in a computer science department). This is related to the second most-identified theme, that there are challenges in changing or adding to the physics curriculum. These respondents believed that the curriculum has a lot of material to cover as is, and some were already engaging in changes. Lastly, some respondents do not have experience using data science, and others were unsure how to include it in their courses.
Some other themes include students having little experience programming (N = 9), faculty bandwidth to change the curriculum in their courses (N = 4), and lack of decision-making power to change the curriculum (N = 3).
Industry Survey
25 individuals from 22 different employers responded to this survey. The companies that employed these respondents varied in size, from 1 to 1000s. Most worked in government scientific research (N = 10) or in tech (N = 7). Others worked in for-profit research (general), finance, advertising, health care, and social media, with most working in companies that covered multiple areas. Most (N = 19 or 76%) worked with data scientists with physics degrees.
Figure 3: Results of data skills survey for industry practitioners. Relevant skills are listed on the x-axis, with the different textures showing the proportion of responses received.
Respondents were asked about various technical skills and knowledge they would expect from an entry-level data scientist with a bachelor’s in physics. Most (N = 20 or 80%) would expect such an employee to be programming daily, with slightly fewer (N = 18 or 72%) using various data science software packages (e.g., NumPy harris2020array , SciPy 2020SciPy-NMeth , Pandas reback2020pandas ) daily. Almost all respondents indicated that such employees use Python (N = 22, 88%), though many (N = 16, 64%) use C/C++ and Matlab (N = 15, 60%). In figure 3, over half of the industry respondents indicated that the following were essential for an entry-level data scientist job: Data Manipulation (e.g., transforming data, gaining insights from data); Experimental design and scientific method (e.g., designing experiments, data collection); Statistics (e.g., model selection and uncertainty, A/B testing); and Data Visualization and Communication (e.g., plotting data, technical writing).
Additionally, some respondents left further clarifying comments. They emphasized that tools (e.g., programming languages) are less important than understanding how to make useful plots or understanding the bigger goals. Good collaboration and communication skills and best practices for software development and error analysis were also mentioned.
Challenges
As demonstrated in the survey, there are also several challenges in incorporating data science into the undergraduate physics curriculum. First, while some physicists have the necessary background to teach data science and machine learning, this is not true for all. Many physicists who teach undergraduate courses are not familiar enough with data science topics to teach them, and preparing such new content is extremely time-consuming for them.
A second challenge that presents itself when trying to incorporate data science into the physics classroom is finding a place for it in the normal course structure. Especially for courses taught at an institution for many semesters, a well-established flow of topics leaves little room for new material to be added. Additionally, many universities lack the resources to offer a new class specifically on data science topics, and many students lack the flexibility in their schedules to take an additional course.
The final challenge we wish to discuss is the required background students must have outside the physics curriculum. In order to fully understand much of machine learning, a higher level of mathematics, statistics, and programming is needed. Some of these skills may be picked up in mathematics or computer science courses but are not typically taught in the physics classroom. Additionally, to understand introductory data science, students will need to be familiar with basic programming. While knowledge of programming is becoming more common among physics students (even becoming a requirement in some departments), this is not necessarily true for all universities and thus hinders incorporating data science into the physics classroom when students do not have the necessary background knowledge.
II Data Science Education Community of Practice (DSECOP)
The mission of DSECOP dsecop_web is to support physics educators who wish to integrate data science into their existing courses by providing resources as well as a community of like-minded educators for support. The DSECOP project came to life in 2022 following the awarding of the competitive Innovation Fund Award from the American Physical Society to some leaders of the Topical Group of Data Science (GDS) aps_gds .
One of the resources provided by the DSECOP organization is a collection of modules created by the group’s fellows who are Ph.D. students and recent Ph.D. graduates. Each module introduces a specific topic in the realm of data science but is applied to a physics concept taught in the traditional physics major. The goal is that these modules are detailed enough to provide instructors and students with a good understanding of how data science can be used to solve a problem in physics but also short enough to be incorporated into a pre-existing course without significant alterations. These modules consist of a lesson plan, exercises, and solutions for the instructor. We adopted this approach because not all universities will have the resources to develop a full course on data science. This also allows for the content to be introduced gradually into the curriculum by individual professors rather than requiring a new course approval that has to replace a well-established physics course. It also presents alternative materials to the students for further learning.
Our fellows present an initial idea for a module and, after refining it with discussion, send it out for outside feedback from faculty. They then develop a unit, which is later sent out for testing by other faculty. We also run an annual workshop, where the fellows and faculty present their work, followed by a discussion. We run a webinar series to present different approaches to teaching data science. These have ranged from ethics to careers in industry.
II.1 DSECOP Modules
Data science (and its sub-field machine learning) requires many levels of knowledge that build on each other before students are capable of tackling more complicated problems. To streamline the usage of modules, we establish workflows and conventions for both module creators and faculty members. We assume basic Python knowledge and standardize the list of libraries such as Matplotlib Hunter2007matplotlib , NumPy harris2020array , SciPy 2020SciPy-NMeth , Pandas reback2020pandas , TensorFlow tensorflow2015-whitepaper . Setting up Python environments across different machines can be challenging. Additionally, not all students might have access to GPUs, which are integral for modern data science tasks. To address these issues, all the modules can be run on the Google Colab platform GoogleColaboratory , allowing students to run the modules in their browsers on a cloud GPU without any extra setup.
The DSECOP modules span a wide range of concepts over data science and machine learning; thus, we have sought to order them regarding the level of data science knowledge. We introduce the DSECOP numbering system, where each module is assigned a label of the form DSECOP XXX, where XXX is a three-digit module ID. A module ID starting with a 1 means that the module contains data science concepts appropriate for new and beginner programmers. 200-level modules denote modules that have intermediate-level machine learning concepts, and thus, the students should have experience with programming and some of the concepts listed in the module’s prerequisites to complete the modules. 300-level modules are the most advanced modules we currently offer, which cover a variety of advanced machine-learning concepts. A roadmap of the DSECOP modules and their suggested pre-requisite modules is shown in Figure 4. This flowchart breaks the modules up into those which contain general data science topics, and those which contain machine learning topics.
Figure 4: This flowchart shows the relationships between the modules and their topic (general data science or machine learning). The arrows depict suggested prerequisites to cover the relevant data science and machine learning topics, starting from the most basic module DSECOP 101: Introduction to Data Science Libraries. Modules that cover data science and machine learning at the introductory level are shown with solid borders and are numbered DSECOP 1XX. Modules with intermediate topics have dashed borders and are numbered DSECOP 2XX and modules that cover advanced topics in data science and machine learning have dotted borders and are numbered DSECOP 3XX.
Our most introductory module is “DSECOP 101: Introduction to Data Science Libraries” (link), which only assumes that students have a basic knowledge of Python and introduces the students to performing a simple data analysis with the Python libraries Pandas, Seaborn, and Matplotlib. This module is connected to all other modules in Figure 4 as the data science and machine learning concepts become more advanced. For example, one of the more advanced modules offered, “DSECOP 235: Schrödinger’s Equation and Neural Networks” (described in detail later in this paper), uses a complicated neural network set up to solve Schrödinger’s equation. If an instructor wanted to use this module in a course but feared their students’ knowledge of neural networks may not be strong enough, the instructor could also assign one or two of the prerequisite modules, such as “DSECOP: Differential Equations and Neural Networks” or “DSECOP 201: Introduction to Deep Learning”.
At the time of this paper’s publication, we had 21 posted modules spanning many common courses in the undergraduate physics curriculum. Table 2 breaks down the current modules by their physics content and divides them among the common courses in a physics curriculum. Note that the modules “DSECOP 201: Introduction to Deep Learning” and “DSECOP 210: Introduction to Random Forests” do not contain any physics content and thus are applicable in any physics course with the proper computational background.
General Physics (Lecture and Lab Waves and Optics Classical Mechanics
DSECOP 101: Intro to DS Libraries
DSECOP 110: Intro to Histograms
textbfDSECOP 120: Exploratory Data Analysis
DSECOP 140: ML Workflow
DSCECOP 150: Intro to Classification
DSECOP 160: Outlier Analysis
DSECOP 170: Error Reduction
DSECOP 220: Spectral Clustering
DSECOP 230: Diff Eqs + NN
DSECOP 290: Time Series + RNN
Electricity and Magnetism Quantum Mechanics Statistical Thermodynamics
DSECOP 190: Data Science with Griffiths
DSECOP 250: Automatic Differentiation
DSECOP 235: Schrodinger + NN
DSECOP 236: Schrodinger + GP
DSECOP 320: Monte Carlo + NN
Advanced Lab Particle, Nuclear, or Medical Physics Any Course
DSECOP 120: Exploratory Data Analysis
DSECOP 240: Symbolic Regression
DSECOP 280: Automated Video Analysis
DSECOP 330: SVD
DSECOP 101: Intro to DS Libraries
DSECOP 110: Intro to Histograms
DSECOP 340: NMR + NN
DSECOP 102: The Professor’s Module
DSECOP 201: Intro to Deep Learning
DSECOP 210: Intro to Random Forests
Table 2: An overview of courses in the undergraduate physics curriculum and existing modules for each course.
The next two sections of this paper examine two of these DSECOP modules in more detail. “DSECOP 110: Introduction to Data Processing with Histograms” is one of our introductory modules giving an overview of creating meaningful graphs with histograms and “DSECOP 235: Scrödinger’s Equation and Neural Networks” is a more advanced module which seeks to solve Schödinger’s equation with physics-informed neural networks.
II.1.1 Introduction to Data Processing with Histograms
This module is designed to teach students the steps involved in a data analysis pipeline. It focuses on key concepts such as processing datasets, creating histograms, curve fitting, and determining the goodness of fit for a chosen model. It covers the practical use of tools such as NumPy and SciPy for creating histograms and curve fitting.
The module comprises in-class notebooks, a quiz, a homework assignment, and relevant datasets. Four notebooks guide students through the process of analyzing a dataset, including the creation of histograms and curve fitting. A short quiz is also included to assess students’ understanding. The homework assignment then requires students to undertake a practical data analysis task drawing from the material presented in the notebooks. Several datasets are provided for this purpose, along with Jupyter notebooks for instructors to re-generate the datasets if desired. It is estimated that the completion of each notebook will take approximately 60 minutes. The quiz is designed to take around 30 minutes, and the homework assignment is anticipated to require less than 90 minutes of commitment.
The module explores the application of histograms as a tool for data visualization, where students are expected to create their own histogram from a raw dataset and employ built-in functions. Students are introduced to the concept of interpreting histograms as probability distributions. This concept is exemplified using a toy “Uranium-241” dataset, where students are expected to import the data, create a histogram, and subsequently normalize these histograms. The module is designed to be a part of an undergraduate laboratory course or a particle physics course.
Students are expected to achieve several physics and data science competencies after completing this module. From a physics perspective, they should demonstrate the capability to use histograms to visualize large datasets, evaluate how effectively a chosen model aligns with a dataset, and construct their own data analysis pipelines. On the data science front, the students should become adept at utilizing built-in functions from numpy and scipy to produce histograms from a dataset, using these libraries to curve fit a distribution, and even writing their own code for these tasks, if desired. A bonus advanced topic available to students is the utilization of the Kolmogorov-Smirnov goodness-of-fit test for histograms, providing an additional level of depth for those wishing to further their understanding. Pre-requisites for this module include familiarity with Python, numpy, and matplotlib, which are crucial to understanding and completing the exercises and tasks included in the module.
II.1.2 Learning the Schrödinger Equation
This module aims to introduce students to deep learning as applied to the Time-Dependent Schrödinger Equation (TDSE), a fundamental equation in quantum mechanics (QM), through a combination of analytical as well as computational (ML) solutions. The module highlights Physics-Informed Neural Networks (PINNs), a class of machine learning algorithms that facilitate the solution of partial differential equations by integrating constraints from partial differential equations within deep learning models. The module is designed for an undergraduate QM course, and it guides students through solving the TDSE with a focus on the Quantum Harmonic Oscillator (QHO). The only prerequisite is an introductory understanding of Python and the basics covered in the initial weeks of a QM course.
The module is divided into three submodules. The first is a gentle, interactive introduction to deep learning, where students learn to create neural network pipelines through a simple example involving RGB color mixing. The second part consists of an article that offers an overview of various ML concepts with an emphasis on physics applications. The third submodule provides an interactive notebook for solving the TDSE for a 1D QHO system, consisting of two states in superposition as they evolve in time.
Supplementing the main content within the notebooks are hands-on lessons, exercises, homework and suggested projects. These materials allow students to apply what they have learned and deepen their understanding of the subject. There are code cells that students are expected to fill themselves, references for further study, take-home assignments, and suggestions for topics that build on the material and can be used as research projects. The estimated time for students to complete the module is about six hours, in addition to two hours taught by a professor in a classroom setting.
The module can be accessed via Google Colab, with GPU support for faster performance, or through a local environment using Conda. Detailed instructions are provided for both methods.
The “Learning the Schrödinger Equation” module blends traditional quantum mechanics with modern machine learning techniques. Through a structured sequence of submodules, students are exposed to the TDSE and are then led to explore it through hands-on applications of deep learning. Students are encouraged to extend this module by exploring other quantum systems, such as particle-in-a-box or double-well potential, and to apply the techniques from this module for solving partial differential equations encountered in other physical systems.
III Conclusions and Outlook
This paper highlights the imperative integration of data science into the undergraduate physics curriculum, emphasizing the growing interplay between these two fields. With a growing number of physicists employing data science tools in their research, the necessity for this incorporation becomes ever more apparent. Physics methodology is also being used to understand and extend data science tools. Physicists entering the workforce in the industry, national labs, or academia will be well served by having exposure to these methods. This mirrors the requirement for physicists to possess proficiency in mathematical methods and computer programming.
We introduced DSECOP as a means to address this need. Educators interested in including data science in their courses are welcome to join the community. Educators can choose units based on their teaching subjects and their comfort with the methodology. In the future, we plan to hold workshops at national conferences where instructors can have more direct interaction with the graduate fellows who develop our modules.
Acknowledgements
The DSECOP project was funded and actualized with the support of the 2021 American Physical Society Innovation Fund award. This project would not have been possible without the support of Dr. Marilena Longobardi and Dr. Wolfgang Losert. We acknowledge useful discussions with Dr. Jackson and Dr. Craig from EP3, Dr. Robert Hilborn from AAPT, and Dr. Talitha Washington from Atlanta University Center. We extend our acknowledgment to our DSECOP Fellows for their invaluable contributions, without which this project would not have achieved its success; 2022 fellows, listed in alphabetical order: Sebastian Atalla, Fatemeh Bagheri, Julie Butler, Cunwei Fan, Radha Mastandrea, and Karan Shah; and 2023 fellows, also in alphabetical order: Julie Butler, Ashley Dale, Joseph Dominicus Lap, Richard Harry, Connor Robertson, and Karan Shah. Support for WR’s participation in this education and outreach activity was provided by the CHRNS project under a partnership between the NIST and the NSF (DMR-2010792). Any mention of commercial products within this paper is for information only; it does not imply recommendation or endorsement by NIST.
Conflict of Interest
The authors have no conflicts to disclose.